Grade 11Algebra

logx (x^3-x^2-2x)

harshit choudhary
10 Years agoGrade 11

1 Answer

BALAJI ANDALAMALA
10 Years ago
log_x(x^3-x^2-2x)=log_x(x(x^2-x-2))
=log_xx+log_x(x^2-x-2)
=1+log_x((x+1)(x-2)) = 1+log_x(x+1)+log_x(x-2)